# HyReaL: Clustering Attributed Graph via Hyper-Complex Space Representation Learning

### v2

@@ 1 Introduction
  • Quaternion is novelly introduced to appropriately address both the OS and OD effects. For OD, FVP feature amplification and QGE feature coupling learning enhancement collaboratively emphasize the graph attributes. For OS, the ‘4$\times$×DoF’ of quaternion transformation ensures the learning ability of the model, making it unnecessary to tack too many GCN layers.
@@ 4.1 Experimental Settings
Six evaluation metrics are utilized. Three external metrics Zhou et al. (2022a): Clustering Accuracy (ACC), Normalized Mutual Information (NMI), and Average Rand Index (ARI), which evaluate performance according to the data labels, are in the intervals $[0,1]$, $[0,1]$, and $[-1,1]$, respectively. Three internal metrics: Silhouette Coefficient (SC) Rousseeuw (1987), Davies-Bouldin Index (DBI) Davies and Bouldin (1979), and Calinski-Harabasz Index (CHI) Caliński and Harabasz (1974), that do not rely on the labels, are in the intervals $[-1,1]$, $[0,+\infty)$, and $[0,+\infty)$. All these metrics are commonly used by most of the compared state-of-the-art methods, and except for DBI, a higher value indicates a better clustering performance.
@@ A.2 Counterparts and HyReaL Settings
There are some hyper-parameters and settings of our method, i.e., the layer number, pre-training learning rate, pre-training iteration number, learning rate, iteration number, model regularization trade-off $\alpha$, and representation embedding loss trade-off $\beta$. We set Adam optimizer during experiments. The activation function of the graph encoder is ReLU for each layer. $\mathcal{L}_{reg}$ in the loss is the regularization of model, and L1 regularization is utilized. In the pre-training process, the hyper-parameter $\beta$ is set to $0.0001$. For ten datasets, the neuron number of layers, the pre-training learning rate, pre-training iteration number, and iteration number are set to $[512,256,128]$, $10^{-4}$, $10$, and $4$, respectively. The learning rate is set to $0.00001$ for CITESEER, $0.0002$ for ACM, WIKI, and AMAP, $0.0004$ for DBLP, FILM, CORNELL, CORA, WISC, and, UAT. The $\alpha$ is set to $10^{-4}$ for CORNELL, $2\times 10^{-4}$ and UAT, $5\times 10^{-4}$ for CITESEER, FILM, and WISC, $10^{-5}$ for CORA and DBLP, $10^{-6}$ for ACM, WIKI, AMAP. The $\beta$ is set to $2^{-10}$ for WIKI, DBLP, CORNELL, CORA, WISC, $2^{-12}$ for ACM, CITESEER, $2^{-30}$ for UAT, 00 for AMAP.
@@ A.3 Description of Validity Metrics
Here, $TP$ and $FP$ respectively denote the number of true positive pairs and true negative pairs, and $C_{n}^{2}$ is the number of possible object pairs.
@@ A.4 Settings of the Wilcoxon Signed-ranks Test
Table A.2: The Wilcoxon signed rank test with 95% confidence interval. The symbols “$+$” and “--” indicate the rejection and acceptance of the null hypothesis.
@@
K-Means $+$ $+$ $+$
Spectral-C $+$ $+$ $+$
GAE $+$ $+$ $+$
VGAE $+$ $+$ $+$
ARGAE $+$ $+$ $+$
ARVGAE $+$ $+$ $+$
CONVERT -- $+$ $+$
CCGC $+$ $+$ $+$
DFCN $+$ $+$ $+$
DAEGC $+$ $+$ $+$
EGAE $+$ $+$ $+$
@@ B.3 Computational Complexity Analysis
The time complexity of the proposed HyReaL model is $\mathcal{O}(T[nd\hat{d}+n^{2}\hat{d}^{2}])$. We analyze it below. The training process of the model is composed of three parts: (1) quaternion projection, (2) quaternion graph convolution, and (3) graph reconstruction. For the quaternion projection, the dimensions of the input and projected features of each projector are $d$ and $\hat{d}$, respectively. Since four MLP layers are paralleled to project the attribute values of the $n$ nodes, the time complexity is thus $\mathcal{O}(4nd\hat{d})$. For the quaternion graph convolution, the feature quaternion in size $n\times(4\times\hat{d})$ will be processed by $l$ stacked quaternion graph encoders. The parameters $\mathbf{W}^{Q}_{l}$ of each encoder are with the same scale as the feature quaternion. Hence, the time complexity of quaternion graph convolution is $\mathcal{O}(l(n4\hat{d})^{2})$. For the graph reconstruction, the inner product is conducted on the matrix $\mathbf{\Gamma}$ with size $n\times\hat{d}$, which consumes $\mathcal{O}(n^{2}\hat{d})$. Assume the training of HyReaL iterates $T$ times, the overall time complexity is $\mathcal{O}(T[4nd\hat{d}+l(n4\hat{d})^{2}+n^{2}\hat{d}])$. By omitting the small constants and the terms with lower magnitude, the final complexity is nearly $\mathcal{O}(T[nd\hat{d}+n^{2}\hat{d}^{2}])$.
@@ C.6 Bonferroni-Dunn Test of Comparison Experiment
The Bonferroni-Dunn test is used to compare an algorithm with the remaining $k-1$ counterparts. It involves comparing the differences in average ranks of various methods with a certain threshold value called Critical Difference (CD). The CD is defined as:
@@
where $q_{\lambda}$ is critical values for the BD test, $p$ is the number of compared methods, and $N$ is the number of dataset. If the rank difference between the two methods is higher than the CD, it indicates that the method with the higher average rank is statistically superior to the one with the lower average rank. Conversely, if the difference is lower than the CD, it suggests that there is no significant performance difference between the two methods.
@@
where the $q_{0.10}$ and $q_{0.05}$ of ten classifiers are 2.539 and 2.773 according to Table 5(b) in reference Demsar [2006], the number of datasets $N$ is 10, and the number of compared methods $p$ is 10. Overall, it can be observed that HyReaL performs significantly better than the seven methods, as shown in Figure A.1.
@@ D.1 Detailed Remark of Learning Ability
[…]
@@ D.2 Proof of Degree of Freedom
Rigorous proof of the ‘4$\times$×DoF’ is provided below. We follow the parameter initialization method in Parcollet et al. [2019] and initial our quaternion graph encoders. In order to simplify the mathematical expression and analyses, especially when dealing with operations involving complex numbers or quaternions, we introduce the polar coordinate form to represent the weight of quaternion. A generated quaternion weight $w$ from a weight matrix $\mathbf{W}^{Q}$ has a polar form defined as:
@@
with $q_{img}^{\triangleleft}=0+x\mathbf{i}+y\mathbf{j}+z\mathbf{k}$ a purely imaginary and normalized quaternion, and the angle $\theta$ is randomly generated in the interval $[-\pi,\pi]$. The imaginary components $x\mathbf{i}$, $y\mathbf{j}$, and $z\mathbf{k}$ are sampled from an uniform distribution in $[0,1]$ to obtian $q_{imag}$. The parameter $\varphi$ is a random number generated with respect to well-known initialization criterions (such as Glorot Glorot and Bengio [2010] or He algorithms He et al. [2015]). Therefore, $w$ can be computed following:
@@
$[\mathbb{E}(|\mathbf{W^{Q}}|)]^{2}$ equals to 0 since the weight distribution is symmetric around 0. Nonetheless, the value of $\operatorname{Var}(\mathbf{W^{Q}})=\mathbb{E}\left(|\mathbf{W^{Q}}|^{2}\right)$ is not trivial in the case of quaternion-valued matrices. Indeed, $\mathbf{W^{Q}}$ follows a Chi-distribution with four degrees of freedom (DoFs). Chi-distribution is often used to describe the distribution mode of the modulus length or Euclidean distance of multi-dimensional vectors. Thus, $\mathbb{E}\left(|\mathbf{W^{Q}}|^{2}\right)$ is expressed and computed as follows:
@@
[…]

### v3

@@ 1 Introduction
[…]
@@ 4.1 Experimental Settings
Six evaluation metrics are utilized. Three external metrics Zhou et al. (2022a): Clustering Accuracy (ACC), Normalized Mutual Information (NMI), and Average Rand Index (ARI), which evaluate performance according to the data labels, are in the intervals $[0,1]$, $[0,1]$, and [1,1][-1,1], respectively. Three internal metrics: Silhouette Coefficient (SC) Rousseeuw (1987), Davies-Bouldin Index (DBI) Davies and Bouldin (1979), and Calinski-Harabasz Index (CHI) Caliński and Harabasz (1974), that do not rely on the labels, are in the intervals [1,1][-1,1], $0,+\infty)$, and $0,+\infty)$. All these metrics are commonly used by most of the compared state-of-the-art methods, and except for DBI, a higher value indicates a better clustering performance.
@@ A.2 Counterparts and HyReaL Settings
There are some hyper-parameters and settings of our method, i.e., the layer number, pre-training learning rate, pre-training iteration number, learning rate, iteration number, model regularization trade-off $\alpha$, and representation embedding loss trade-off $\beta$. We set Adam optimizer during experiments. The activation function of the graph encoder is ReLU for each layer. $\mathcal{L}_{reg}$ in the loss is the regularization of model, and L1 regularization is utilized. In the pre-training process, the hyper-parameter $\beta$ is set to $0.0001$. For ten datasets, the neuron number of layers, the pre-training learning rate, pre-training iteration number, and iteration number are set to $[512,256,128]$, $^{-4}, 1010$, and 44, respectively. The learning rate is set to $0.00001$ for CITESEER, $0.0002$ for ACM, WIKI, and AMAP, $0.0004$ for DBLP, FILM, CORNELL, CORA, WISC, and, UAT. The $\alpha$ is set to $10^{-4}$ for CORNELL, $2\times 10^{-4}$ and UAT, $5\times 10^{-4}$ for CITESEER, FILM, and WISC, $10^{-5}$ for CORA and DBLP, $10^{-6}$ for ACM, WIKI, AMAP. The $\beta$ is set to $2^{-10}$ for WIKI, DBLP, CORNELL, CORA, WISC, $2^{-12}$ for ACM, CITESEER, $2^{-30}$ for UAT, 0 for AMAP.
@@ A.3 Description of Validity Metrics
Here, TPTP and FPFP respectively denote the number of true positive pairs and true negative pairs, and $C_{n}^{2}$ is the number of possible object pairs.
@@ A.4 Settings of the Wilcoxon Signed-ranks Test
Table A.2: The Wilcoxon signed rank test with 95% confidence interval. The symbols “++” and “-” indicate the rejection and acceptance of the null hypothesis.
@@
K-Means ++ ++ ++
Spectral-C ++ ++ ++
GAE ++ ++ ++
VGAE ++ ++ ++
ARGAE ++ ++ ++
ARVGAE ++ ++ ++
CONVERT - ++ ++
CCGC ++ ++ ++
DFCN ++ ++ ++
DAEGC ++ ++ ++
EGAE ++ ++ ++
@@ B.3 Computational Complexity Analysis
[…]
@@ C.6 Bonferroni-Dunn Test of Comparison Experiment
The Bonferroni-Dunn test is used to compare an algorithm with the remaining k1k-1 counterparts. It involves comparing the differences in average ranks of various methods with a certain threshold value called Critical Difference (CD). The CD is defined as:
@@
where $q_{\lambda}$ is critical values for the BD test, $p$ is the number of compared methods, and NN is the number of dataset. If the rank difference between the two methods is higher than the CD, it indicates that the method with the higher average rank is statistically superior to the one with the lower average rank. Conversely, if the difference is lower than the CD, it suggests that there is no significant performance difference between the two methods.
@@
where the $q_{0.10}$ and $q_{0.05}$ of ten classifiers are 2.539 and 2.773 according to Table 5(b) in reference Demsar [2006], the number of datasets NN is 10, and the number of compared methods $p$ is 10. Overall, it can be observed that HyReaL performs significantly better than the seven methods, as shown in Figure A.1.
@@ D.1 Detailed Remark of Learning Ability
$\begin{aligned} \mathbf{F}^{Q}&=\mathbf{F}\otimes\mathbf{W}^{Q}\\ &=\mathbf{W}^{Q}_{r}\mathbf{F}_{r}-\mathbf{W}^{Q}_{x}\mathbf{F}_{x}-\mathbf{W}^{Q}_{y}\mathbf{F}_{y}-\mathbf{W}^{Q}_{z}\mathbf{F}_{z}\\ &+\mathbf{W}^{Q}_{x}\mathbf{F}_{r}+\mathbf{W}^{Q}_{r}\mathbf{F}_{x}-\mathbf{W}^{Q}_{z}\mathbf{F}_{y}+\mathbf{W}^{Q}_{y}\mathbf{F}_{z}\\ &+\mathbf{W}^{Q}_{y}\mathbf{F}_{r}+\mathbf{W}^{Q}_{z}\mathbf{F}_{x}+\mathbf{W}^{Q}_{r}\mathbf{F}_{y}-\mathbf{W}^{Q}_{x}\mathbf{F}_{z}\\ &+\mathbf{W}^{Q}_{z}\mathbf{F}_{r}-\mathbf{W}^{Q}_{y}\mathbf{F}_{x}+\mathbf{W}^{Q}_{x}\mathbf{F}_{y}+\mathbf{W}^{Q}_{r}\mathbf{F}_{z}\end{aligned},$ (9)
@@ D.2 Proof of Degree of Freedom
[…]
@@
with $q_{img}^{\triangleleft}=0+x\mathbf{i}+y\mathbf{j}+z\mathbf{k}$ a purely imaginary and normalized quaternion, and the angle $\theta$ is randomly generated in the interval $\pi,\pi]$. The imaginary components $x\mathbf{i}$, $y\mathbf{j}$, and $z\mathbf{k}$ are sampled from an uniform distribution in $[0,1]$ to obtian $q_{imag}$. The parameter $\varphi$ is a random number generated with respect to well-known initialization criterions (such as Glorot Glorot and Bengio [2010] or He algorithms He et al. [2015]). Therefore, $w$ can be computed following:
@@
$\mathbb{E}(|\mathbf{W^{Q}}|)]^{2}$ equals to 0 since the weight distribution is symmetric around 0. Nonetheless, the value of $\operatorname{Var}(\mathbf{W^{Q}})=\mathbb{E}\left(|\mathbf{W^{Q}}|^{2}\right)$ is not trivial in the case of quaternion-valued matrices. Indeed, $\mathbf{W^{Q}}$ follows a Chi-distribution with four degrees of freedom (DoFs). Chi-distribution is often used to describe the distribution mode of the modulus length or Euclidean distance of multi-dimensional vectors. Thus, $\mathbb{E}\left(|\mathbf{W^{Q}}|^{2}\right)$ is expressed and computed as follows:
@@
With $P(x)$ is polynomial and has a limit to $\mathbf{+\infty}$}. The second term is calculated in a same way (integration by parts) and $\mathbb{E}(|\mathbf{W}^{Q}|^{2})$ becomes from Eq. (27):